# A Global Minimum Tax for Large Firms Only:

### v2

@@ A.1. Proof of Proposition 1
A.2. Proofs of Propositions 3 and 4
@@
A.3. Regime-switching points
@@ A.4. Proof of Proposition 5
A.5. Decentralized decision of tax havens
In the text, we have assumed that tax havens decide whether to set a single GMT rate or split their tax rate collectively. We here examine an alternative decentralized decision process. More specifically, each single tax haven decide whether to split or not given the other havens’ decisions.
Suppose that $t_{M}\in[t_{M}^{0},t_{M}^{2}]$, and there are $H_{M}$ havens that commit to a single GMT and the remaining $H_{S}=H-H_{M}$ havens that split their tax rate. When the non-haven sets a single non-GMT rate and a single haven sets a single GMT rate, the equilibrium tax rates are
which is the counterpart of (A.1). Let $G_{h}^{1}(H_{M})$ be the haven’s equilibrium payoff.
When the non-haven sets a single non-GMT rate and a single haven splits its tax rates, the equilibrium tax rates are
which is the counterpart of (A.2). Let $G_{h}^{2}(H_{M})$ be the haven’s equilibrium payoff.
We then ask if a single haven has an incentive to choose a single GMT rate given that all the other $H-1$ havens split their tax rates. The single haven has an unilateral incentive to do so if the following holds:
$\Delta G_{h}^{12}(H_{M}=1)$ $\equiv G_{h}^{1}(H_{M}=1)-G_{h}^{2}(H_{M}=0)$
$=\frac{(1-\phi)[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta}{\delta H^{2}[\lambda(3+\phi)-(1+\phi)]^{2}[\phi(\lambda-1)(H-1)+\lambda(3H+1)-(H+1)]}\geq 0,$
$Ht_{M}(3\lambda-1)-\delta(\lambda-1)\geq 0\ \ \text{or}\ \ t_{M}\geq\frac{\delta(\lambda-1)}{H(3\lambda-1)}\equiv t_{M}^{0}.$
which is equivalent with $t_{M}\leq t_{M}^{a}\equiv A/B$. Supposing that this inequality holds, the same reasoning applies to all the symmetric havens. This results in the same equilibrium as Regime 1 in the text, where all the havens commit to a single GMT rate and the non-haven sets a single non-GMT rate.
However, for Regime 1 to be sustainable, any single haven must have no incentive to split given that all the other $H_{M}-1$ havens commit to a single GMT rate. The single haven has no unilateral incentive to split its tax rate if the following holds:
$\Delta G_{h}^{12}(H_{M}=H)$ $\equiv G_{h}^{1}(H_{M}=H)-G_{h}^{2}(H_{M}=H-1)$
$=\frac{(1-\phi)[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta^{\prime}}{\delta H(2\lambda-1)[\phi(\lambda-1)+\lambda(4H-1)-(2H-1)]^{2}}\geq 0,$
$D\equiv(6H^{2}-4H+1)\lambda^{2}-(5H^{2}-6H+2)\lambda+(H-1)^{2}-\phi(\lambda-1)^{2}>0,$
and $t_{M}\geq t_{M}^{0}$ holds again. The equivalent condition is $t_{M}\leq t_{M}^{b}\equiv C/D$. We can check $t_{M}^{a}<t_{M}^{b}<t_{M}^{1}<t_{M}^{2}$ and immediately see $\Delta G_{h}^{12}(H_{M}=H)<0$ for $t_{M}\in(t_{M}^{b},t_{M}^{2}]$.
Therefore we can conclude that for $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, all the havens set a single GMT rate and the non-haven sets a single non-GMT rate; for $t_{M}\in(t_{M}^{a},t_{M}^{2}]$, all the havens split its tax rate and the non-haven sets a single non-GMT rate. The former regime exactly corresponds to Regime 1 and the latter one to Regime 2 in the text. The only difference from the text is that the range of Regime 1 under havens’ decentralized decision process is smaller than in the text, implying that the havens’ commitment is less likely. This is intuitive because the commitment of a single haven benefits itself by raising its own tax rate, but the single commitment does not induce the non-haven to increase its tax rate as much as the commitment of many havens.
We can also confirm that in the remaining range of GMT, $t_{M}\in(t_{M}^{2},1]$, a single haven always has an unilateral incentive to split given that all the other havens split, whether the non-haven set a single GMT rate or split. The results of our quantitative exercise in Section 4 remain unchanged because, at $t_{M}=0.15$, the calibrated economy is in Regime 2 anyway.

### v3

@@ A.1. Proof of Proposition 1
A.2. Decentralized decision of tax havens
In our main model, we have assumed that tax havens collectively decide on whether to set a single GMT rate, or to split their tax rate. We here examine an alternative, decentralized decision process in the first stage of our game where each single tax haven decides whether to split or not, given the other havens’ decisions. To simplify this analysis, we assume that countries can split their tax rate only if they are bound by the GMT. This reflects the fact that preferential tax regimes for specific tax bases are not allowed under the OECD’s action plan to fight base erosion.
Suppose then that $t_{M}\in[t_{M}^{0},t_{M}^{2}]$. Out of the total of $H$ tax havens, there are $H_{M}$ havens that commit to a single GMT rate while the remaining $H_{S}=H-H_{M}$ havens split their tax rate. When the non-haven sets a single non-GMT rate and a single haven $h$ sets a uniform GMT rate, the equilibrium tax rates are
which is the counterpart of (A.1). Let $G_{h}^{1}(H_{M})$ be the haven’s equilibrium payoff with commitment.
If the single haven $h$ splits its tax rate instead, whereas the non-haven sets a uniform non-GMT rate, the equilibrium tax rates are
which is the counterpart of (A.2). Let $G_{h}^{2}(H_{M})$ be the haven’s equilibrium payoff under splitting.
We then ask if a single haven has an incentive to choose a uniform GMT rate given that all the other $H-1$ havens split their tax rates. A single haven $h$ has a unilateral incentive to do so, if the following holds:
$\Delta G_{h}^{12}(H_{M}=1)$ $\equiv G_{h}^{1}(H_{M}=1)-G_{h}^{2}(H_{M}=0)$
$=\frac{(1-\phi)[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta}{\delta H^{2}[\lambda(3+\phi)-(1+\phi)]^{2}[\phi(\lambda-1)(H-1)+\lambda(3H+1)-(H+1)]}\geq 0,$ (A.9)
$\text{and}\qquad Ht_{M}(3\lambda-1)-\delta(\lambda-1)\geq 0\ \ \text{or}\ \ t_{M}\geq\frac{\delta(\lambda-1)}{H(3\lambda-1)}\equiv t_{M}^{0}.$
It is straightforward to show that $\Theta>0$ holds at $t_{M}^{0}$ and hence, by continuity, also for $t_{M}$ marginally above $t_{M}^{0}$. Therefore, there must be a positive range of $t_{M}$, with $t_{M}^{0}<t_{M}\leq t_{M}^{a}\equiv A/B$ for which the inequality (A.9) holds for a haven $h$. By symmetry, this argument applies to all havens.
For a Regime 1 equilibrium with commitment to occur, a further condition is that no single haven $h$ must have an incentive to split, given that all the other $H_{M}-1$ havens commit to a single GMT rate. A single haven $h$ has no unilateral incentive to split its tax rate, if the following holds:
$\Delta G_{h}^{12}(H_{M}=H)$ $\equiv G_{h}^{1}(H_{M}=H)-G_{h}^{2}(H_{M}=H-1)$
$=\frac{(1-\phi)[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta^{\prime}}{\delta H(2\lambda-1)[\phi(\lambda-1)+\lambda(4H-1)-(2H-1)]^{2}}\geq 0,$ (A.10)
$D\equiv(6H^{2}-4H+1)\lambda^{2}-(5H^{2}-6H+2)\lambda+(H-1)^{2}-\phi(\lambda-1)^{2}>0.$
and $t_{M}\geq t_{M}^{0}$ holds again. The equivalent condition is $t_{M}\leq t_{M}^{b}\equiv C/D$. We can immediately see $\Delta G_{h}^{12}(H_{M}=H)<0$ for $t_{M}\in(t_{M}^{b},t_{M}^{2}]$ and check $t_{M}^{a}<t_{M}^{b}<t_{M}^{1}<t_{M}^{2}$:
$t_{M}^{b}-t_{M}^{a}=\frac{\delta(H-1)(1-\phi)^{2}(\lambda-1)^{4}}{\Theta^{\prime\prime}[(6H^{2}-4H+1)\lambda^{2}-(5H^{2}-6H+2)\lambda+(H-1)^{2}-\phi(\lambda-1)^{2}]}>0,$
$t_{M}^{1}-t_{M}^{b}=\frac{\delta(H-1)(2\lambda-1)(\lambda-1)^{2}}{\Theta^{\prime\prime}[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}>0,$
$\Theta^{\prime\prime}\equiv 3(3H-1)\lambda^{2}-2(3H-2)\lambda+H-1+\phi(\lambda-1)[\lambda(3H-5)-(H-3)]>0.$
We can conclude that for $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, all the havens set a single GMT rate and the non-haven sets a single non-GMT rate; for $t_{M}\in(t_{M}^{a},t_{M}^{2}]$, all the havens split their tax rate and the non-haven sets a single non-GMT rate. The former regime corresponds to Regime 1 and the latter one to Regime 2 in the text. The only difference is that the range of Regime 1 is smaller when havens make their first stage decision in a decentralized way, implying that the havens’ commitment is less likely. This is intuitive because the commitment of a single haven does not induce the non-haven to increase its tax rate by as much as the simultaneous commitment of many havens. Finally, in the remaining range of the GMT, $t_{M}\in(t_{M}^{2},1]$, the difference between collective vs. decentralized decisions of havens plays no role any more. In either setting, the tax havens always have an incentive to split their tax rate, whether the non-haven sets a single GMT rate (Regime 3), or splits its tax rate as well (Regime 4).
A.3. Proofs of Propositions 3 and 4
@@
A.4. Regime-switching points